Question 7 071 pts Using the following information, lead times are one week for items P and B two weeks for items A and K, and three weeks for items C and D. An order for P is scheduled to be shipped at the start of week 13. What is the latest week that the order can be started and still be ready to ship on time? B(5) C(3) K(2) K(3) D(3) Week 10 Week 8 TOSHIBA
Answer #1

Correct Answer:

Week 7

Working note:

D-A-P = 3+2+1 = 6 weeks

K-A-P = 2+2+1 = 5 weeks

K-B-P = 2+ 1+ 1 = 4 weeks

K-C-P = 2+ 3+ 1 = 6 weeks

Here, the longest time is 6 week and it is the critical time. So, it will take 6 week to complete the order. Since, the shipment will start from week 13.

So, latest time = 13-6 = week 7